# On the existence of unbiased resilient estimators in discrete quantum systems

### v2

@@ Abstract
Cramér-Rao constitutes a crucial lower bound for the mean squared error of an estimator in frequentist parameter estimation, albeit paradoxically demanding highly accurate prior knowledge of the parameter to be estimated. Indeed, this information is needed to construct the optimal unbiased estimator, which is highly dependent on the parameter. Conversely, Bhattacharyya bounds result in a more resilient estimation about prior accuracy by imposing additional constraints on the estimator. Initially, we conduct a quantitative comparison of the performance between Cramér-Rao and Bhattacharyya bounds when faced with less-than-ideal prior knowledge of the parameter. Furthermore, we demonstrate that the $n^{th}$order classical and quantum Bhattacharyya bounds cannot be computed –given the absence of estimators satisfying the constraints– under specific conditions tied to the dimension $m$ of the discrete system. Intriguingly, for a system with the same dimension $m$, the maximum non-trivial order $n$ is $m-1$ in the classical case, while in the quantum realm, it extends to $m(m+1)/2-1$. Consequently, for a given system dimension, one can construct estimators in quantum systems that exhibit increased robustness to prior ignorance.
@@ I Introduction
Quantum metrology studies the limits in precision by estimating physical parameters encoded in quantum systems. Quantum properties such as entanglement and squeezing are used to surpass classical estimation strategies [1, 2, 3]. Applications like quantum phase estimation [4, 5, 6, 7], quantum-enhanced position and velocity estimation [8, 9, 10], quantum illumination [11, 12, 13], quantum thermometry [14, 15, 16, 17], and channel discrimination [18] among others have boosted the field. Parameter estimation gives the theoretical framework to establish the optimal performance of particular estimation protocols and its main figure of merit is the Cramér-Rao bound (CRB) [19, 20].
Paradoxically, to achieve the ultimate precision established by the CRB an infinite number of measurements have to be performed or the prior value of the parameter has to be known. Both scenarios are unlikely or experimentally unfeasible. In the finite sample regime, CRB is frequently unattainable [21, 22, 23]. Several bounds as Hammersley–Chapman–Robbins bound [24], the family of Bhattacharyya bounds (BhBB) [25] or Barankin [26, 27] bounds, and more generally Abel bounds [28], which include the previous bounds as special cases, were introduced. These bounds succeed in establishing a sensible limit for the achievable precision [29], especially in noisy estimation problems [27].
Recently, there has been interest in developing tighter bounds for quantum parameter estimation tasks. Gessner et al. [30] derived lower bounds on the variance of estimators in quantum metrology by increasing the unbiasedness constraints of the estimator, i.e., in the same spirit as their classical analogs were obtained. They presented a hierarchy of increasingly tight bounds that include the quantum CRB (QCRB) [31] at the lowest order and quantum Barankin, Bhattacharyya, and Abel bounds as other particular cases.
In this article, we study the classical and the quantum Bhattacharyya bounds (QBhBB). We compare the performance of these bounds with the CRB when the prior knowledge of the parameter is not sharp deriving a lower bound for the mean squared error (MSE). We illustrate this result for the Mach-Zehnder interferometer by comparing the actual MSE of the estimator with BhB and CRB and we see a clear advantage when using the BhB. Finally, we address the impossibility of computing classical BhB under certain conditions when dealing with discrete probability distributions. We prove the necessary and sufficient conditions for this bound to exist and we show how these conditions get relaxed for the QBhB.
This article is organized as follows: in sec. II, we introduce estimation theory, present the BhB and motivate them using the Mach-Zehnder interferometer setting. In sec. III, we study the cases when the BhB cannot be computed, i.e. the bound is divergent for both, classical probability distributions and quantum systems. Finally, in sec. IV, the conclusions are presented.
@@ II Estimation theory
To extract information from a physical system measurements have to be performed. The measurement outcome is modeled as a random variable following a probability distribution that depends on the state of the system. The information is encoded in the random variable and estimation theory is the branch of statistics that deals with the optimal protocols and the ultimate limit for the information extraction. The CRB establishes a limit on the precision while estimating a magnitude or a parameter. Consider that the outcome of the experiment, denoted by the random variable $X$, follows the family of probability distribution $P_{\theta}(x)$, that is parameterized by the parameter $\theta$ which is being estimated. An estimator $\tilde{\Theta}(x)$ is a function of the random measurement outcome that estimates the value of the parameter $\theta$. The performance of an estimator is characterized by its bias $b(\theta)$, mean squared error (MSE) and variance $(\Delta\tilde{\Theta})^{2}$,
[…]
$(\Delta\tilde{\Theta}_{\theta})^{2}\geq\frac{1}{F_{C}},$ (4)
@@
[…]
@@
where $X_{+}$ is the space of events $x$ with nonzero probability $P_{\theta}(x)>0$.
@@
The CRB is always saturable in the asymptotic limit or when the true value of $\theta$ is known [32]. This is not the case in realistic scenarios. In quantum parameter estimation [1], a family of quantum states $\rho_{\theta}$ parametrized by a parameter $\theta$ plus a particular POVM measurement $\{\Pi_{x}\}$ substitute the family of probability distributions, $P_{\theta}(x)$. The lower bound on the MSE of the estimator is now determined by the quantum Cramér-Rao bound (QCRB) [1, 31, 14], its version for locally unbiased estimators reads
@@
From Eq.(7) it is deduced that the SLD is a hermitian operator and its eigenvectors can form an orthonormal basis. In the derivation of the QCRB, a general POVM $\{\Pi_{x}\}$ was considered but it was shown that a projection on the basis of eigenvectors of $L$ is the optimal measurement that saturates the QCRB [14]. As in the classical estimation theory, the QCRB is always saturable in the asymptotic regime. Furthermore, it has recently been shown to be non-saturable within a Bayesian perspective [33, 23].
II.1 Bhattacharyya bounds and their motivation
Bhattacharyya bounds were introduced in the context of classical parameter estimation problems to give a lower bound for the variance of the estimate [25]. The BhB gives a tighter bound than the CRB by including higher-order derivatives of the probability distribution in the calculation. It is shown that BhB converge to the variance of the best unbiased estimator –the estimator that is unbiased for any point in the region of the allowed values of the parameter–, when the sampling distribution is a member of an exponential family of distributions [34]. The $n^{th}$order BhB bound gives the lowest value of the variance for an estimator that fulfills $\frac{d^{i}b(\theta)}{d\theta^{i}}|_{\theta=\theta_{0}}=0$ for $1\leq i\leq n$. This estimator is unbiased in a larger region than the one obtained for the CRB, as will be shown later, and its variance is closer to the real variance of a reasonable parameter estimation problem, namely where the real value of the parameter is unknown, and is given by [30],
$\bm{\lambda}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix},\>\>\>\>\>\>\>\bm{C}_{kl}=\int_{x\in X_{+}}\frac{\partial^{k}P_{\theta_{0}}(x)\partial^{l}P_{\theta_{0}}(x)}{P_{\theta_{0}}(x)},$ (13)
where
We would like to point out the similarities between the $n^{th}$order BhB bound and the multiparameter Cramér-Rao bound (MCRB) [35] in a multiparameter estimation problem with $n$ parameters. The computational effort to compute both bounds is the same and the BhB can be codified as a particular case of the CRB in a multiparameter estimation setting. The BhB can be extended to quantum estimation theory [30] in a coherent way that includes them as a specific case of a larger set of bounds. Doing so, the QBhB state that the variance of any unbiased estimator with null $n$ first bias derivatives satisfies
@@ II.1 Bhattacharyya bounds and their motivation
where $L_{l}$ is the generalization of the symmetric logarithmic derivative that satisfies
@@
II.2 An example: Bhattacharyya bound in Mach-Zehnder interferometer
Historically more attention has been paid to the Barakin bounds [26, 27]. The Bhattacharyya bounds have not been computed in any physical problem before. Here, we compute BhB and the estimators achieving the bound for the paradigmatic Mach-Zehnder interferometer. We show that when the prior is not sharp, and we move away from the point for which the unbiased estimators were constructed, the BhB bound offers an improvement to the achievable precision.
$\varphi_{a}$$\varphi_{b}$$a$$b$$a^{\prime}$$b^{\prime}$|$\ket{r}$|$\ket{r}$$n_{a}$$n_{b}$
Figure 1: Mach-Zehnder interferometer setup
II.2.1 Mach-Zehnder
In the Mach-Zehnder configuration, depicted in Fig. (1), two modes of the electromagnetic field interfere on a balanced beamsplitter, the output beams acquire a relative phase $\theta=\varphi_{a}-\varphi_{b}$, and interfere again on the second beamsplitter. Finally, the photon numbers $n_{a}$ and $n_{b}$ are measured at the output ports. We consider the input state |$\ket{r,r}$ where both input modes are in the Fock state with $r$ photons. It has been shown that this state achieves the Heisenberg limit [36] while estimating the relative phase $\theta$. The evolution induced by the interferometer is given by the unitary $U=e^{\frac{\theta}{2}(a^{\dagger}b-b^{\dagger}a)}$ [6]. The probability distribution $P_{\theta}(2q)$ is then
$P_{\theta}(2q)=J^{2}_{q}(r\theta),$ (18)
[…]
II.3 Analytical motivation
In this section, we derive a lower bound for the MSE and show that it depends on the bias. We first define $\hat{\Theta}_{\theta_{0}}(x)$ as an estimator unbiased at $\theta_{0}$. We compute the MSE of this estimator in the point $\theta^{\prime}$
[…]
We now use the Cauchy-Schwarz inequality to find a lower bound for Eq. (19). Doing so, one gets
[…]
Making the choice $f(x)=(\theta^{\prime}-\tilde{\Theta}_{\theta_{0}})\sqrt{P_{\theta^{\prime}}(x)}$ and $g(x)=\sqrt{P_{\theta^{\prime}}(x)}$, we obtain
[…]
The first integral is the square of the bias, the second is the MSE, and the last integral is just the unity. So we can conclude that
$b^{2}(\theta^{\prime})\leq\text{MSE}(\theta^{\prime}).$ (22)
The squared bias gives a lower bound for the MSE. Since the derivatives of the bias for the BhB estimator are zero, this estimator will have a smaller lower bound for the MSE for an interval $\theta_{0}-\epsilon<\theta^{\prime}<\theta_{0}+\epsilon$. Unfortunately, the saturability of (22) highly depends on the particular probability distribution and it is not possible in general. Nevertheless, in Appendix (A) we prove that if the probability distribution satisfies certain conditions there always exists an interval where the $n^{th}$order BhB performs at least as well as the CRB.
@@ III Conditions of the existence of estimators
Up to this point, we presented the usefulness of the BhB for experimentally friendly scenarios. However, one cannot always construct estimators that fulfill the conditions needed to lower bound their MSE by the BhB. This is true especially if one deals with discrete probability distributions. In this section, we give the conditions for the existence of the estimators for finite support probability distributions. We conclude that in some cases, higher-order BhB might not be computable or irrelevant. The largest BhB computable – and that gives us new information– is related to the dimension of the support of the probability distribution. We also show that the BhB diverges if and only if no estimator is satisfying the unbiasedness conditions.
We present our results for both classical and quantum parameter estimation settings. The problem of quantum parameter estimation encompasses two key optimization challenges: (i) the optimization across all conceivable observables or, more broadly, the optimization across all feasible POVMs that can be executed; and (ii) the optimization across all potential estimators $\tilde{\Theta}$ that can be derived from the outcomes of measurements. While optimization (ii) is addressed by the classical Cramér-Rao bound, optimization (i) introduces a distinctly quantum dimension to the problem. We will demonstrate the implications of this extra freedom to select a POVM.

### v3

@@ Abstract
The Cramér-Rao bound serves as a crucial lower limit for the mean squared error of an estimator in frequentist parameter estimation. Paradoxically, it requires highly accurate prior knowledge of the estimated parameter for constructing the optimal unbiased estimator. In contrast, Bhattacharyya bounds offer a more robust estimation framework with respect to prior accuracy by introducing additional constraints on the estimator. In this work, we examine divergences that arise in the computation of these bounds and establish the conditions under which they remain valid. Notably, we show that when the number of constraints exceeds the number of measurement outcomes, an estimator with finite variance typically does not exist. Furthermore, we systematically investigate the properties of these bounds using paradigmatic examples, comparing them to the Cramér-Rao and Bayesian approaches.
preprint: APS/123-QED
@@ I Introduction
Quantum metrology investigates the fundamental limits of precision in estimating physical parameters encoded in quantum systems. Taking advantage of inherently quantum features such as entanglement and squeezing, it enables precision levels beyond those attainable by classical estimation strategies [1, 2, 3]. Applications span quantum phase estimation [4, 5, 6, 7], quantum-enhanced position and velocity estimation [8, 9, 10], quantum illumination [11, 12, 13], quantum thermometry [14, 15, 16, 17], and quantum channel discrimination [18], driving substantial progress in the field. The theoretical framework of parameter estimation provides benchmarks for the optimal performance of quantum-enhanced estimation protocols.
Parameter estimation has traditionally been approached from two main perspectives: frequentist and Bayesian [19]. These frameworks, based on different interpretations of probability, yield distinct insights into parameter estimates and their uncertainties. Although frequentist and Bayesian approaches converge in the asymptotic limit of large measurement numbers, this has often led to the misconception that they are interchangeable in quantum metrology. In reality, they reflect fundamentally different paradigms. In the frequentist framework, the Cramér-Rao bound (CRB) [20, 21] is the main figure of merit, providing a lower bound on the mean squared error of estimators. However, achieving the precision dictated by the CRB paradoxically requires either a vast number of measurements or highly accurate prior knowledge of the parameter—both of which are often impractical or experimentally infeasible. In finite-sample regimes, the CRB frequently becomes unattainable [22, 23, 24]. To address these limitations, alternative bounds have been developed, including the Hammersley–Chapman–Robbins bound [25], the Bhattacharyya bounds (BhBB) [26], the Barankin bounds [27, 28], and the more general Abel bounds [29]. These bounds are particularly valuable in noisy estimation scenarios, providing more realistic precision limits [28].
Recent efforts have aimed to derive tighter bounds for quantum parameter estimation. For example, Gessner et al. [30] proposed a hierarchy of lower bounds on estimator variance by imposing additional unbiasedness constraints. This hierarchy includes the quantum CRB (QCRB) [31] as its lowest-order bound, while higher-order cases include quantum analogs of the Barankin, Bhattacharyya, and Abel bounds.
In this work, we analyze both classical and quantum Bhattacharyya bounds (QBhBB) and assess their performance relative to the CRB and Bayesian approaches, particularly in scenarios with uncertain prior knowledge of the parameter. We first illustrate these bounds with two paradigmatic cases, namely, the estimation of the variance of a Gaussian distribution, and the analysis of the performance of a Mach-Zehnder interferometer setup. Our findings reveal that the BhBB-inspired strategy can provide a significant advantage in specific regimes. Furthermore, we investigate the conditions under which the classical Bhattacharyya bound becomes non-computable, deriving the necessary and sufficient criteria for the existence of these bounds. Notably, we demonstrate that the quantum Bhattacharyya bound has more relaxed conditions for computability compared to its classical counterpart.
This article is organized as follows: In Sec. II, we introduce the Bhattacharyya bounds, discuss why they are tighter bounds, and highlight scenarios in which they are advantageous. In Sec. III, we explore cases where the Bhattacharyya bound diverges and becomes non-computable, and we establish the necessary conditions for the existence of these bounds in both classical probability distributions and quantum systems. In Sec. IV, we analyze the Mach-Zehnder interferometer, comparing the precision scaling achieved by the Cramér-Rao bound and the Bhattacharyya bounds. Finally, we summarize our findings and present our conclusions.
@@ II Estimation theory
To extract information from a physical system, measurements must be performed. The measurement outcome is modeled as a random variable that follows a probability distribution dependent on the state of the system. The information is encoded in this random variable, and estimation theory is the branch of statistics that addresses optimal protocols and the ultimate limits for information extraction. The Cramér-Rao bound (CRB) establishes a limit on the precision of estimating a quantity or parameter. Consider that the outcome of the experiment, denoted by the random variable $X$, follows the family of probability distributions $P_{\theta}(x)$, parameterized by the parameter $\theta$, which is being estimated. An estimator $\tilde{\Theta}(x)$ is a function of the random measurement outcome that provides an estimate of the parameter $\theta$. The performance of an estimator is characterized by its bias $b(\theta)$, mean squared error (MSE), and variance $(\Delta\tilde{\Theta})^{2}$.
$b(\theta)$ $=\sum_{x}(\theta-\tilde{\Theta}(x))P_{\theta}(x)\>,$ (1)
$(\Delta\tilde{\Theta})^{2}$ $=\sum_{x}(\langle\tilde{\Theta}\rangle_{\theta}-\tilde{\Theta}(x))^{2}P_{\theta}(x)\>,$ (2)
MSE $=\sum_{x}(\theta-\tilde{\Theta}(x))^{2}P_{\theta}(x)\>.$ (3)
The mean value is defined as $\langle f\rangle_{\theta}=\sum_{x}f(x)P_{\theta}(x)\>$. An estimator $\tilde{\Theta}_{\theta_{0}}(x)$ is locally unbiased on the point $\theta_{0}$ when $\langle\tilde{\Theta}_{\theta_{0}}\rangle_{\theta_{0}}=\theta_{0}$ and $\frac{db(\theta)}{d\theta}|_{\theta=\theta_{0}}=0$. Note that for unbiased estimators MSE and the variance coincide. Unbiased Cramér-Rao, usually named as the Cramér-Rao bound (CRB), states that for any unbiased estimator
$(\Delta\tilde{\Theta}_{\theta_{0}})^{2}\geq\frac{1}{F_{C}},$ (4)
@@
$F_{C}=\sum_{x\in X_{+}}\left(\frac{\partial P_{\theta}}{\partial\theta}\right)^{2}\frac{1}{P_{\theta}(x)}\>,$ (5)
@@
where $X_{+}$ is the space of events $x$ with nonzero probability $P_{\theta}(x)>0$. Previous definitions hold for discrete probability distributions but can be generalized for continuous distributions replacing sums with integrals.
The Cramér-Rao bound (CRB) can always be saturated in the asymptotic limit or by using a specific local estimator that depends on $\theta$ [32]. However, in realistic scenarios where $\theta$ is unknown, the CRB does not serve as a tight bound. Bayesian parameter estimation, in contrast, is a global approach, as Bayesian-inspired estimators do not rely on local knowledge of $\theta$. Unfortunately, Bayesian methods are often challenging to compute [33].
II.1 Bhattacharyya bound
Bhattacharyya bounds (BhB) were introduced in the context of classical frequentist parameter estimation problems to provide a lower bound for the variance of an estimate [26]. The BhB gives a tighter bound than the Cramér-Rao bound (CRB) by incorporating higher-order derivatives of the probability distribution in the calculation. It is shown that the BhB converges to the variance of the best unbiased estimator—the estimator that is unbiased for any point in the region of the allowed values of the parameter—when the sampling distribution belongs to an exponential family of distributions [34].
The $n^{\text{th}}$-order BhB gives the lowest value of the variance for an estimator that satisfies
$\frac{d^{i}b(\theta)}{d\theta^{i}}\bigg{|}_{\theta=\theta_{0}}=0,\quad\text{for }1\leq i\leq n,$
where $b(\theta)$ denotes the bias. This estimator is unbiased over a larger region than that obtained using the CRB, as will be shown later, and its variance is closer to the actual variance in realistic parameter estimation problems—namely, cases where the true value of the parameter is unknown. The BhB is given by [30],
$\bm{\lambda}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix},\>\>\>\>\>\>\>\bm{C}_{kl}=\sum_{x\in X_{+}}\frac{\partial^{k}P_{\theta_{0}}(x)\partial^{l}P_{\theta_{0}}(x)}{P_{\theta_{0}}(x)},$ (7)
where
The bound is in principle saturable by an estimator that satisfies,
$\bm{\lambda}^{\top}\bm{C}^{-1}\bm{g}(x)=P_{\theta_{0}}(x)\left(\tilde{\Theta}_{\theta_{0}}(x)-\theta_{0}\right)$ (8)
and $\bm{g}(x)^{T}=(\partial P_{\theta_{0}}(x),\partial^{2}P_{\theta_{0}}(x),...,\partial^{n}P_{\theta_{0}}(x))$. We would like to point out the similarities between the $n^{th}$order BhB bound and the multiparameter Cramér-Rao bound (MCRB) [35] in a multiparameter estimation problem with $n$ parameters. The BhB can be codified as a particular case of the CRB in a multiparameter estimation setting.
For a limited number of measurements, Bayesian estimation techniques are often considered more appropriate than frequentist methods (see, e.g., Refs. [36, 22] for discussion). In the Bayesian framework, one assumes a prior distribution $p_{\text{prior}}(\sigma)$ that reflects knowledge about the parameter $\sigma$ before any measurements are made. The prior is then updated according to Bayes’ rule, resulting in the Bayesian posterior distribution $p_{\text{post}}(\sigma|x)$ after observing the measurement outcome $x$, given by
$p_{\text{post}}(\sigma|x)=\frac{p(x|\sigma)\,p_{\text{prior}}(\sigma)}{\int\mathrm{d}\theta\,p(x|\sigma)\,p_{\text{prior}}(\sigma)}.$ (9)
The posterior distribution is used to induce an estimation for $\sigma$, given the outcome $x$.
II.2 Application: Gaussian Distribution
We consider the problem of estimating the variance $\sigma$ of a zero mean Gaussian probability distribution,
$P_{\sigma}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{\frac{-x^{2}}{2\sigma^{2}}}.$ (10)
(a) Mean squared error of different strategies optimized for $\sigma_{0}=1$ when the true value of the parameter is $\sigma$.
(b) Bias given by different strategies optimized for $\sigma_{0}=1$ when the true value of the parameter is $\sigma$.
Figure 1: Comparison of the MSE (upper figure) and the bias (lower figure) for a Gaussian distribution using Bayesian, CR, BhB2 and BhB5 strategies.
We compare the mean squared error and bias of several estimators saturating Bhattacharyya bounds for the point $\sigma_{0}=1$ and given by (8) with the estimator saturating Cramér-Rao in the same point and a Bayesian estimator. The Bayesian estimator considered is the value of $\sigma$ that maximizes the posterior, that is,
$\tilde{\sigma}(x)=\arg\max_{\sigma}p_{\text{post}}(\sigma|x).$ (11)
In the Bayesian approach, the information about $\theta$ provided by the posterior probability always depends on the chosen prior distribution $p_{\text{prior}}(\theta)$. If prior knowledge about $\theta$ is available, it can be incorporated by selecting an appropriate prior distribution. However, in the absence of prior information, determining a ”noninformative” prior becomes challenging. A flat prior, $p_{\text{prior}}(\theta)=\text{const}$, may appear noninformative, but it is not invariant under reparameterizations, meaning it could encode unintended information for different functions of $\theta$.
To address this, we adopt Jeffreys prior, which is given by $p_{\text{prior}}(\theta)\propto\sqrt{F_{C}}$, where $F_{C}$ is the classical Fisher information. Jeffreys prior is particularly useful as it remains invariant under reparameterization, ensuring consistency across different representations of the parameter. In Fig. [1], we depict the mean squared error (MSE) and the bias of the different estimators. As expected, the Bayesian approach gives a lower MSE over a large range of values of $\sigma$, but its bias is the largest. On the other hand, the estimator satisfying the order-5 Bhattacharyya bound (BhB) has the smallest bias in the interval considered, but its MSE increases rapidly for large values of $\sigma$. The most suitable estimator is the one saturating the order-2 BhB, as its MSE remains more constant than $\text{MSE}_{\text{CR}}$, with a bias close to zero.
II.3 Quantum bounds
@@
In quantum parameter estimation [1], a family of quantum states $\rho_{\theta}$, parameterized by a parameter $\theta$, along with a particular POVM measurement $\{\Pi_{x}\}$, replaces the family of probability distributions $P_{\theta}(x)$. The lower bound on the mean squared error (MSE) of the estimator is now determined by the quantum Cramér-Rao bound (QCRB) [1, 31, 14]. Its version for locally unbiased estimators is given by
@@
From Eq. (13) it is deduced that the SLD is a hermitian operator and its eigenvectors form an orthonormal basis. In the derivation of the QCRB, a general POVM $\{\Pi_{x}\}$ was considered but it was shown that a projection on the basis of eigenvectors of $L$ is the optimal measurement that saturates the QCRB [14]. As in the classical estimation theory, the QCRB is always saturable in the asymptotic regime. Furthermore, it has recently been shown to be non-saturable within a Bayesian perspective [37, 24]. The BhB can be extended to quantum estimation theory [30] in a coherent way that includes them as a specific case of a larger set of bounds. Doing so, the QBhB state that the variance of any unbiased estimator with null $n$ first bias derivatives satisfies
@@
where $L_{l}$ is the generalization of the symmetric logarithmic derivative that obeys
@@
The bound in Eq. (14) is saturated by the estimator that satisfies
$\bm{\lambda}^{\top}\bm{Q}^{-1}\bm{g}(x)=P_{\theta_{0}}(x)\left(\tilde{\Theta}_{\theta_{0}}(x)-\theta_{0}\right).$ (17)
@@ III Conditions of the existence of estimators
Up to this point, we have presented the usefulness of the Bhattacharyya bound (BhB) for experimentally friendly scenarios. However, one cannot always construct estimators that fulfill the conditions required to lower bound their mean squared error (MSE) by the BhB. This is particularly true when dealing with discrete probability distributions.
In this section, we provide the conditions for the existence of such estimators for finite-support probability distributions. We conclude that, in some cases, higher-order BhB may not be computable or relevant. The largest computable BhB that provides new information is related to the dimension of the support of the probability distribution. Furthermore, we show that the BhB diverges if and only if no estimator satisfies the unbiasedness conditions.
In quantum parameter estimation, given a quantum state $\rho_{\theta}$, the estimation process involves two steps: performing a measurement using a POVM and applying classical post-processing to the measurement outcomes. The efficiency of the estimation protocol depends on both the choice of the POVM and the method used to process the measurement data. Consequently, optimal quantum parameter estimation involves tackling two key optimization tasks: (i) optimizing over all possible observables or, more generally, all feasible POVMs that can be implemented, and (ii) optimizing over all potential estimators $\tilde{\Theta}$ derived from the measurement outcomes.
While the classical Cramér-Rao bound addresses the second optimization, the first introduces a uniquely quantum aspect to the problem. In this work, we highlight the implications of this additional freedom in selecting the POVM.